# Neural Risk Limiting Dispatch in Power Networks: Formulation and Generalization Guarantees

### v1

@@ III Neural Risk Limiting Dispatch
Motivated by the above challenges, we propose the neural RLD. In this approach, the hypothesis class $\mathcal{H}$ is represented by L2 regularized neural networks with a finite number of parameters. Consequently, problem (4) is redefined as a finite-dimensional neural network training task with a specialized training loss. To support this approach, we provide a generalization performance guarantee based on statistical learning tools including uniform convergence and Rademacher complexity, which theoretically bounds the optimality gap of our method with high probability. [^2] It is impossible to obtain performance guarantees here with probability one as there is always a non-zero probability (as long as the training set is finite) to obtain a training set for which the empirical mean in (4) is not a good representation of the population mean in (3). This guarantee demonstrates that as more historical data points are utilized, the performance of our neural RLD closely approximates that of the optimal hypothesis in the same hypothesis class. In this section, we first describe the overall neural RLD procedure. Following this, we delve into the training process of the neural RLD. Finally, we establish the generalization performance guarantee.
@@ III-A Neural network for data-driven RLD
The proposed neural RLD chooses L2-regularized neural networks as the hypothesis class due to their excellent approximation capability [30]. By using $k\in\mathcal{K}$ and $j\in\mathcal{J}^{k}$ with $K:=|\mathcal{K}|$ and $J^{k}:=|\mathcal{J}^{k}|$ to index the hidden layers and the neurons in each, this hypothesis class can be expressed as:
@@
where $\mathbf{h}:\mathbb{R}^{p}\mapsto\mathbb{R}^{N_{\mathrm{B}}}$ represents a neural network in this hypothesis class; $\mathbf{W}^{k}\in\mathbb{R}^{J^{k+1}\times J^{k}}$ is the weights of the $k$-th layer and $\mathbf{w}_{j}^{k}\in\mathbb{R}^{J^{k}}$ is the transpose of its $j$-th row. Symbol $\phi(\cdot)$ denotes the activation function. This paper chooses ReLU as the activation function. It operates on each individual element of the input, ensuring that the dimensions of both input and output remain identical. The second constraint in (5) regularize the class of neural networks by limiting the L2 norm of the weights, where $W^{\max}$ is the maximum norm of weights allowed. If we further define a new loss function $\ell$, also referred as to the RLD loss, as follows:
@@
then problem (4) can be interpreted as a neural network training task. This task uses $\mathcal{S}=\{(\mathbf{x}^{(m)},\mathbf{d}^{(m)}),\forall m\in\mathcal{M}\}$ as the training set, and leverages $\ell$ as the training loss, as follows:
@@ III-B Neural network training
where $\mathbf{u}\in\mathbb{R}^{{N_{\mathrm{B}}}}$ is the output of the hypothesis $\mathbf{h}$, and we have $\nabla_{\mathbf{w}_{j}^{k}}\ell\in\mathbb{R}^{J^{k}}$, $\nabla_{\mathbf{u}}Q\in\mathbb{R}^{{N_{\mathrm{B}}}}$, and $\nabla_{\mathbf{w}_{j}^{k}}\mathbf{u}\in\mathbb{R}^{{N_{\mathrm{B}}}\times J^{k}}$. The gradient $\nabla_{\mathbf{w}_{j}^{k}}\mathbf{u}$ can be easily obtained using automatic differentiation in modern deep learning frameworks like Pytorch. However, the gradient $\nabla_{\mathbf{u}}Q$ is hard to calculate because $Q(\mathbf{u},\mathbf{d})$ involves an optimization problem (II-A). To address this issue, we leverage the strong duality to explicitly express $\nabla_{\mathbf{u}}Q$, as described in the following proposition.
Proposition 1 (Expression of the gradient).
The gradient $\nabla_{\mathbf{u}}Q$ can be expressed as:
@@
Then, we can introduce the SGD to train the neural network. The whole training process is summarized in Algorithm 1. Here, we use $N^{\mathrm{ep}}$ and $\mathcal{B}$ to denote the maximum epoch number and the sample index of a batch, respectively.
@@ 4            ${\mathbf{u}}^{(m)}\leftarrow\mathbf{h}({\mathbf{x}}^{(m)}),\forall m\in\mathcal{B}$;
5             Solve the dual of (II-A) to get $(\bm{\lambda}^{\star})^{(m)},\forall{m}\in\mathcal{B}$;
@@
Algorithm 1 Neural network training
@@ III-C Generalization performance guarantee
This section provides a theoretical generalization performance guarantee, which bounds the optimality gap of the proposed neural RLD with a large probability. Specifically, we first define the generalization cost of a hypothesis $\mathbf{h}$, denoted by $L(\mathbf{h})$, as follows:
@@
As introduced in Section III-A, we train a L2-regularized neural network $\widehat{\mathbf{h}}$ to act as our hypothesis. In order to calculate the excess cost of $\widehat{\mathbf{h}}$, two mild assumptions are in force:
@@
Theorem 1 (PAC bound of excess cost).
@@
$\Delta L(\widehat{\mathbf{h}})\leq\frac{4(2{W^{\max}})^{K-\frac{1}{2}}C_{\ell}X^{\mathrm{max}}+\sqrt{2\ln(2/\delta)}}{\sqrt{M}}.$ (12)
@@
By (11), the PAC bound provided by Theorem 1 can be regarded as the expected optimality gap of the hypothesis $\widehat{\mathbf{h}}$. Thus, we also refer to Theorem 1 as the generalization performance guarantee of hypothesis $\widehat{\mathbf{h}}$. This PAC bound decreases to zero as more samples are used, which indicates that the performance of $\widehat{\mathbf{h}}$ gradually approaches to the optimal hypothesis $\mathbf{h}^{\star}$ with probability $1-\delta$ even for arbitrarily small $\delta>0$. Hence, we have the following colloary.
Colloary 1 (Asymptotic optimality).
@@ IV-A1 Test systems
TABLE I: Descriptions of our test cases and structures of the corresponding neural networks
Cases Line number Load number Generator number Structures of of neural networks
@@ IV-A2 Benchmarks
  • Two-step predict-then-optimize paradigm used in [32] (termed “Two-step”). It first follows [14] and trains a mixture density network to learn the distribution of net demands. According to our test, this model shows high accuracy, e.g., the relative errors for predicting the mean and covariance are smaller than 3%. Then, the scenario-based method is employed to solve problem (II-A), in which 30 discrete scenarios are generated to approximate the predicted distribution.
  • Imitation learning used in [21] (termed “Imitation”). This benchmark trains a neural network using the MSE loss to directly learn the mapping from auxiliary information $\mathbf{x}$ to the dispatch decision $\mathbf{u}$. The training labels, i.e., historical optimal decisions, are generated by solving the deterministic dispatch problems given the realizations of $\mathbf{d}$ in the past delivery intervals.
@@ IV-A3 Data generation
$\mathbf{d}=\bm{\Omega}\left(\mathbf{x}*(\bm{1}+0.15\cdot\bm{\pi})\right),$ (13)
where $\bm{\Omega}\in\mathbb{R}^{{N_{\mathrm{B}}}\times p}$ is a fixed coefficient matrix; *** represents the element-wise multiplication. A random variable $\bm{\pi}\in\mathbb{R}^{p}$ is introduced to simulate the stochastic characteristic of net demands. Here, we assume that $\bm{\pi}$ follows the standard Gaussian distribution. 4,000 samples are treated as the training set, while the rest ones are used as the testing set to verify the suboptimality of different methods.
IV-A4 Structures of neural networks
The neural networks used in all methods are composed of fully connected layers. Their structures are summarized in the column “Structures of neural networks” in Table I, where every number in brackets represents the neuron number in a hidden layer. For example, (5, 5, 5) represents the corresponding neural network has three hidden layers, each with five neurons. Note the neural networks maintain the same structures among all methods.
@@ IV-B1 Convergence
Fig. 2 demonstrates the training losses of the neural networks in each test case. Note all the training losses are standardized to [0,1] for a fair comparison. In every case, the proposed method can converge in only 100 epochs. In contrast, both the benchmarks may not converge even after 1,000 epochs. According to Proposition 1, the gradient of the RLD loss is well-defined, so the proposed model can be efficiently trained by the SGD. In benchmark Imitation, the neural network is trained using the MSE to learn the mapping from auxiliary information $\mathbf{x}$ to the optimal decision $\mathbf{u}^{\star}$. Since the RLD problem (1) is not strictly convex, this mapping may be discontinuous [17], which is hard to learn. Benchmark Two-step trains a mixture density network to learn the distribution of net demands for a given $\mathbf{x}$ [14]. This mixture density network needs to predict the means and covariances for multiple Gaussian components, while a set of parameters need to be learned for each Gaussian component. Thus, it has more learnable parameters than the common neuron network even if they have the same hidden layers and neurons, leading to lower training efficiency.
@@
Figure 2: Standardized training losses of the neural networks trained by different methods. Note the horizontal axis is logarithmic.
@@ IV-B2 Suboptimality
where $\widehat{\mathbf{u}}$ represents the dispatch decision given by each method. The numerator represents the excess generation cost between $\widehat{\mathbf{u}}$ and the optimal decision ${\mathbf{u}^{\star}}$, while the denominator is the optimal objective of (II-A). We calculate this normalized suboptimality for every data point in the testing set. The results are given in Fig. 3. In all test cases, the normalized suboptimality of the proposed method is significant smaller than those of the two benchmarks. For instance, the average suboptimality of the proposed method is only 1.46% in the 300-bus test case, while this value is 5.54% and 3.07% in benchmarks Imitation and Two-step, respectively. As aforementioned, the RLD loss can more accurately reflect the suboptimality of decisions compared to the MSE loss used in benchmark Imitation. Moreover, the proposed neural network is an end-to-end optimization method and can bypass the “mismatch” issue of benchmarks Two-step. Thus, it exhibits the lowest suboptimality among all methods.
@@ IV-B3 Computational efficiency
Fig. 4 compares the required time for solving one single instance and 5000 instances of (1) using different methods, where benchmark Two-step is solved in parallel with 30 threads. Since both the proposed neural RLD and benchmark Imitation replace the solving process with the forward pass of neural networks, they can output dispatch decisions in a very short time. On the contrary, benchmark Two-step still needs to solve (1), so it requires much longer time. For example, the solving times required by the proposed method and benchmark Imitation is at least two orders of magnitude smaller than that of benchmark Two-step.

### v2

@@ III Neural Risk Limiting Dispatch
Motivated by the above challenges, we propose the neural RLD. In this approach, the hypothesis class $\mathcal{H}$ is represented by L2 regularized NNs with a finite number of parameters. Consequently, problem (4) is redefined as a finite-dimensional NN training task with a specialized training loss. To support this approach, we provide a generalization performance guarantee based on statistical learning tools including uniform convergence and Rademacher complexity, which theoretically characterizes the suboptimality of our method. [^2] It is impossible to obtain performance guarantees here with probability one, as there is always a non-zero probability (as long as the training set is finite) to obtain a training set for which the empirical mean in (4) is not a good representation of the population mean in (3). This guarantee demonstrates that as more historical data points are utilized, the performance of our neural RLD approaches that of the optimal hypothesis in the same hypothesis class. In this section, we first describe the overall neural RLD procedure. Following this, we delve into the training process of the neural RLD. Finally, we establish the generalization performance guarantee.
@@ III-A Neural network for data-driven RLD
The mapping from $\mathbf{x}$ to $\mathbf{u}^{\star}$ is nonlinear in general because: i) the net demand $\mathbf{d}$ may not have a linear relationship with the day-ahead observable information $\mathbf{x}$, and ii) even if $\mathbf{d}$ is linearly dependent on $\mathbf{x}$, the mapping from $\mathbf{x}$ to $\mathbf{u}^{\star}$ is not necessarily linear because it involves an optimization problem. Hence, we select NNs as the hypothesis class due to their superior approximation capability [30]. To mitigate overfitting, L2 regularization is employed. Compared to L1 regularization, L2 regularization promotes smoother and more stable optimization by penalizing large weights proportionally to their square, which also facilitates faster convergence during training. By using $k\in\mathcal{K}$ and $j\in\mathcal{J}^{k}$ with $K:=|\mathcal{K}|$ and $J^{k}:=|\mathcal{J}^{k}|$ to index the hidden layers and the neurons in each, this hypothesis class can be expressed as: [^3] Unlike many studies that incorporate L2-regularization by adding a penalty term to the training loss, we employ the constraint $\|\mathbf{w}^{k}_{j}\|_{2}\leq W^{\max}$ to simplify the derivation of the proposed performance guarantee. Furthermore, as noted in [31, Section 3.4.1], regularization through penalty terms and constraints is theoretically equivalent.
@@
where $\mathbf{h}:\mathbb{R}^{p}\mapsto\mathbb{R}^{N_{\mathrm{B}}}$ represents a NN in this hypothesis class; $\mathbf{W}^{k}\in\mathbb{R}^{J^{k+1}\times J^{k}}$ is the weights of the $k$-th layer and $\mathbf{w}_{j}^{k}\in\mathbb{R}^{J^{k}}$ is the transpose of its $j$-th row. Symbol $\phi(\cdot)$ denotes the activation function. This paper chooses ReLU as the nonlinear activation function, i.e., $\phi(\mathbf{x})=\max\{\mathbf{x},\bm{0}\}$. Hence, this NN can represent nonlinear decision rules. It operates on each individual element of the input, ensuring that the dimensions of both input and output remain identical. The second constraint in (5) regularizes the class of NNs by limiting the L2 norm of the weights, where $W^{\max}$ is the maximum norm of weights allowed. If we further define a new loss function $\ell$, also referred as to the RLD loss, as
@@
then problem (4) can be interpreted as a NN training task. This task uses $\mathcal{S}=\{(\mathbf{x}^{(m)},\mathbf{d}^{(m)}),\forall m\in\mathcal{M}\}$ as the training set, and leverages $\ell$ as the training loss, as follows:
@@ III-B NN training
where $\mathbf{u}\in\mathbb{R}^{{N_{\mathrm{B}}}}$ is the output of the hypothesis $\mathbf{h}$, and we have $\nabla_{\mathbf{w}_{j}^{k}}\ell\in\mathbb{R}^{J^{k}}$, $\nabla_{\mathbf{u}}Q\in\mathbb{R}^{{N_{\mathrm{B}}}}$, and $\nabla_{\mathbf{w}_{j}^{k}}\mathbf{u}\in\mathbb{R}^{{N_{\mathrm{B}}}\times J^{k}}$. The gradient $\nabla_{\mathbf{w}_{j}^{k}}\mathbf{u}$ can be easily obtained using automatic differentiation in modern deep learning frameworks like Pytorch. However, the gradient $\nabla_{\mathbf{u}}Q$ is hard to calculate because $Q(\mathbf{u},\mathbf{d})$ involves an optimization problem (2). To address this issue, we leverage the strong duality to explicitly express $\nabla_{\mathbf{u}}Q$, as described in the following proposition.
Proposition 1 (Expression of gradients)
The gradient $\nabla_{\mathbf{u}}Q$ can be expressed as [32]:
@@
Then, we can introduce the SGD to train the NN. The whole training process is summarized in Algorithm 1. Here, we use $N^{\mathrm{ep}}$ and $\mathcal{B}$ to denote the maximum epoch number and the sample index of a batch, respectively.
@@ 4            ${\mathbf{u}}^{(m)}\leftarrow\mathbf{h}({\mathbf{x}}^{(m)}),\forall m\in\mathcal{B}$;
5             Solve the dual of (2) to get $(\bm{\lambda}^{\star})^{(m)},\forall{m}\in\mathcal{B}$;
@@
Algorithm 1 NN training
As described in Algorithm 1, the gradient for NN training is computed using the optimal dual variable $\bm{\lambda}^{\star}$. In other words, the training process involves solving a dual optimization problem, which may reduce training efficiency. However, the training is conducted offline and performed only once. During online decision-making, the trained NN can efficiently predict cost-effective dispatch decisions within a very short time. An alternative to solving optimization problems during training is dual learning [33], where a separate NN is trained to directly predict optimal dual variables. While this approach eliminates the need for solving optimization problems during training, it may introduce additional prediction errors that potentially degrade the quality of dispatch decisions. Given the offline nature of our training process and the importance of ensuring high-quality dispatch decisions, we choose to retain the the dual problem solving during NN training.
@@ III-C Generalization performance guarantee
This section provides a theoretical generalization performance guarantee, which characterizes the suboptimality of the proposed neural RLD with a large probability. Specifically, we first define the generalization cost of a hypothesis $\mathbf{h}$, denoted by $L(\mathbf{h})$, as follows:
@@
As introduced in Section III-A, we train a L2-regularized NN $\widehat{\mathbf{h}}$ to act as our hypothesis. In order to calculate the excess cost of $\widehat{\mathbf{h}}$, two mild assumptions are in force:
@@
Theorem 1 (PAC bound of excess cost)
@@
$\Delta L(\widehat{\mathbf{h}})\leq\frac{4(2{W^{\max}})^{K-\frac{1}{2}}C_{\ell}X^{\mathrm{max}}+\sqrt{2\ln(2/\delta)}}{\sqrt{M}},$ (12)
@@
By (11), the PAC bound provided by Theorem 1 can be regarded as the expected optimality gap of the hypothesis $\widehat{\mathbf{h}}$. Thus, we also refer to Theorem 1 as the generalization performance guarantee of hypothesis $\widehat{\mathbf{h}}$. This PAC bound decreases to zero as more samples are used, which indicates that the performance of $\widehat{\mathbf{h}}$ gradually approaches to the optimal hypothesis $\mathbf{h}^{\star}$ with probability $1-\delta$ even for arbitrarily small $\delta>0$. Hence, we have the following corollary.
Corollary 1 (Asymptotic optimality)
@@
Although the neural RLD method and the generalization performance guarantee are established for L2-regularized NNs, they are fully compatible with other types of NNs, including L1-regularized NNs, recurrent neural networks (RNNs), and convolutional neural networks (CNNs). First, regardless of the type of NN employed, the model can always be trained to predict cost-effective first-stage decisions $\mathbf{u}$ based on observable features $\mathbf{x}$. Second, extending the performance guarantee to other types of NNs requires only minor modifications to the derivation process. Detailed descriptions of these modifications are provided in Appendix B.
Note the RLD formulation in this paper is based on the DC power flow model without considering some realistic settings, such as AC power flow formulation, ramping constraints, and production limits, while allowing unrestricted generation reduction. Nevertheless, the proposed neural RLD method can be extended to incorporate these settings. A detailed discussion on these extensions is provided in Appendix C.
@@ IV-A1 Test systems
TABLE I: Descriptions of our test cases and structures of the corresponding NNs
Cases Line number Load number Generator number Structures of of NNs
@@ IV-A2 Benchmarks
  • Two-step predict-then-optimize paradigm used in [35] (termed “Two-step”). It first follows [11] and trains a mixture density network to learn the distribution of net demands. According to our test, this model shows high accuracy, e.g., the relative errors for predicting the mean and covariance are smaller than 3%. Then, the scenario-based method is employed to solve problem (2), in which 30 discrete scenarios are generated to approximate the predicted distribution.
  • Imitation learning used in [36] (termed “Imitation”). This benchmark trains a NN using the MSE loss to directly learn the mapping from auxiliary information $\mathbf{x}$ to the dispatch decision $\mathbf{u}$. The training labels, i.e., historical optimal decisions, are generated by solving the deterministic dispatch problems given the realizations of $\mathbf{d}$ in the past delivery intervals.
Note that benchmark Two-step is a generalized version of the classic RLD method [8]. The classic RLD assumes that future net demand follows a Gaussian distribution [8]. In contrast, benchmark Two-step employs a mixture density network to predict the probability distribution of future net demand, which accommodates both Gaussian and non-Gaussian distributions and thereby enhancing prediction accuracy. As a result, the benchmark is expected to outperform the classic RLD.
@@ IV-A3 Data generation
$\mathbf{d}=\bm{\Omega}\left(\mathbf{x}*(\bm{1}+0.15\cdot\bm{\omega})\right),$ (13)
where $\bm{\Omega}\in\mathbb{R}^{{N_{\mathrm{B}}}\times p}$ is a fixed coefficient matrix; * represents the element-wise multiplication. A random variable $\bm{\omega}\in\mathbb{R}^{p}$ is introduced to simulate the stochastic characteristic of net demands. Here, we assume that $\bm{\omega}$ follows the standard Gaussian distribution. Among the 5,000 samples, 4,000 are used as the training set, while the remaining 1,000 are used as the testing set to evaluate the suboptimality of the different methods in our case study.
IV-A4 Structures of NNs
The NNs used in all methods are composed of fully connected layers. Their structures are summarized in the column “Structures of NNs” in Table I, where every number in brackets represents the neuron number in a hidden layer. For example, (5, 5, 5) represents the corresponding NN has three hidden layers, each with five neurons. For larger power networks, we utilize more neurons in each layer of the NN as we expect a richer hypothesis class is needed. For example, consider the trivial case where $\mathbf{x}=\mathbf{d}$ and $\mathcal{D}(\mathbf{x})$ is a point mass, i.e., the RLD problem reduces to a simple deterministic linear problem. In this case, $\mathbf{u}^{\star}(\mathbf{x})$ is a piecewise function determined by a parametric linear program [37]. The number of pieces corresponds to the number of critical regions, each defined as a set of $\mathbf{x}$ over which the reduced linear problem exhibits the same constraint binding pattern at the solution $\mathbf{u}^{\star}(\mathbf{x})$. Thus, the number of critical regions depends on the number of possible constraint binding patterns. As larger networks have more inequality constraints (e.g., line flow constraints), the resulting mapping $\mathbf{u}^{\star}(\mathbf{x})$ is necessarily more complex and requires more neurons to approximate. In standard machine learning terminology, the process of determining the number of neurons is referred to as model selection and can be done via, e.g., cross-validation.
@@ IV-B1 Convergence
Fig. 3 demonstrates the training losses of the NNs in each test case. Note all the training losses are standardized to [0,1] for a fair comparison. In every case, the proposed method can converge in only 100 epochs. In contrast, both the benchmarks may not converge even after 1,000 epochs.666 The convergence of our proposed method reaches a saturation point (i.e., starts to slow down) with fewer epochs in the first three cases. However, in the 1354-bus case, it appears later than the other two methods. Nevertheless, an earlier saturation point does not inherently indicate better training efficiency or lower suboptimality. For instance, although the saturation point of the proposed method occurs later in the 1354-bus case, its standardized loss nearly stabilizes after 100 epochs. In contrast, the loss curve of benchmark Imitation shows a significant drop around 1000 epochs, while that of benchmark Two-step continues to decrease even beyond 1000 epochs. These variations in the timing of the saturation point across different methods and test cases may result from the complex interplay among various factors, such as the NN’s scale and the definition of the loss function. According to Proposition 1, the gradient of the RLD loss is well-defined, so the proposed model can be efficiently trained by the SGD. In benchmark Imitation, the NN is trained using the MSE to learn the mapping from auxiliary information $\mathbf{x}$ to the optimal decision $\mathbf{u}^{\star}$. Since the RLD problem (1) is not strictly convex, this mapping may be discontinuous [14], which is hard to learn. Benchmark Two-step trains a mixture density network to learn the distribution of net demands for a given $\mathbf{x}$ [11]. This mixture density network needs to predict the means and covariances for multiple Gaussian components, while a set of parameters need to be learned for each Gaussian component. Thus, it has more learnable parameters than the common neuron network even if they have the same hidden layers and neurons, leading to lower training efficiency.
@@
Figure 3: Standardized training losses of the NNs trained by different methods.
@@ IV-B2 Suboptimality
where $\widehat{\mathbf{u}}$ represents the dispatch decision given by each method. The numerator represents the excess generation cost between $\widehat{\mathbf{u}}$ and the optimal decision ${\mathbf{u}^{\star}}$, while the denominator is the optimal objective of (2). We calculate this normalized suboptimality for every data point in the testing set. The results are given in Fig. 4. In all test cases, the normalized suboptimality of the proposed method is significant smaller than that of the two benchmarks. For instance, the average suboptimality of the proposed method is only 1.46% in the 300-bus test case, while this value is 5.54% and 3.07% in benchmarks Imitation and Two-step, respectively. As aforementioned, the RLD loss can more accurately reflect the suboptimality of decisions compared to the MSE loss used in benchmark Imitation. Moreover, the proposed method is an end-to-end approach that merges the prediction and optimization steps. It can bypass the “mismatch” issue encountered in benchmark Two-step, where prediction and optimization are considered separately. Therefore, it demonstrates the lowest suboptimality.
@@ IV-B3 Computational efficiency
Fig. 5 compares the required time for solving one single instance and 5000 instances of (1) using different methods, where benchmark Two-step is solved in parallel with 30 threads. Since both the proposed neural RLD and benchmark Imitation replace the solving process with the forward pass of NNs, they can generate dispatch decisions in a very short time. In contrast, the Two-step benchmark still requires solving a stochastic program after predicting the distribution of net demand, resulting in significantly longer computation time. For instance, the solving times for the proposed method and the Imitation benchmark are at least two orders of magnitude smaller than those of benchmark Two-step.